This is a differential equations course, please answer all the questions.

Find the general solution.

0 10 15. x(t) = 2 1 -2 -(t) 3 2 -1 16. zt) -30 2 1 - 1 -2 -1 1 0 0 (t)
Answer #1

- 11 XCt) 15. 0 0 1 -2 X(+) 3 2 -1 TO obtain general solution of the system(s) to find eigen valme and eigen vector We haveA2 = xdi (A-1)dı = 0 1-1 0 0 2 1-1 - 2 3 2 0 0 에 2 0 -2 ) % ( 68 ) 2 -2 23 234-223 - 0 => 24 = 23 3X1 + 2xy - 22.3 = 0 24 + 2(1-253) 24 = 0 224 + (1-13) -223 - 0 = (1 - iss) My - 2x3 =0 ث and 34 + 2xy + (-1-13) 3 = 0 222 + (-1-253) zz =0 4 % multiply16. XCt) = بل 0 2 T -2 17 0 A @ 2 0 - 2 - 1 Q (A - AI | = 0 -3-7 10 2 = 0 T 0 - 2 + 2 Q => [-3--) ((-1-a)(--)] (-1 – 2C1+x))vector when eigen x==2 24 let e = be the eigen vector x₂ .: (A - 711) = 0 o => /-3+2 2 -1 + 2 оо - 2 0 + 2 2 0 1 12 -1 2 24 +- 2-isz 0 2 1 - ire - 2 -1 l-ir2 -> (-2-isz ) 4 & 2*3 + 2y = 2+ire 2 ir2 x = 0 => ire de -224 - X2 + (1-2 S2) X3 = 0 04 ų i52X(t) = Gredit general solution Guzetzt +que dez *65)*-*-) szi/efitirat (2tisk 2 (1-1/2) E e szi (2+i 12)